# An Efficient Iterative Decoupling Method for Thermo-Poroelasticity Based on a Four-Field Formulation

### v2

@@ 5.2 Tests for the parameters $K$ and $\Theta$
Table 4: Example 1: The errors and convergence orders with decreasing mesh sizes of coupled and iterative decoupling algorithms. $\bm{K}=10^{-6}\bm{I}$
Alg.1, $\Delta t=10^{-3}$
$1/16$ 1.01039e-01 0.00 6.07933e-03 0.00 2.66261e-01 0.00 2.35294e-01 0.00
$1/32$ 2.54193e-02 1.99 1.50919e-03 2.01 1.18339e-01 1.17 1.10485e-01 1.09
$1/64$ 6.36485e-03 2.00 3.76523e-04 2.00 5.62194e-02 1.07 5.42944e-02 1.02
$1/128$ 1.59182e-03 2.00 9.38336e-05 2.00 2.74790e-02 1.03 2.70270e-02 1.01
Alg.2, $\Delta t=5\times 10^{-3}$, iters=5
$1/16$ 1.01040e-01 0.00 6.08908e-03 0.00 2.68502e-01 0.00 2.37202e-01 0.00
$1/32$ 2.54196e-02 1.99 1.51189e-03 2.01 1.18668e-01 1.18 1.10762e-01 1.10
$1/64$ 6.36496e-03 2.00 3.77393e-04 2.00 5.62122e-02 1.08 5.43382e-02 1.03
$1/128$ 1.59189e-03 2.00 9.43212e-05 2.00 2.74213e-02 1.04 2.70369e-02 1.01
Alg.2, $\Delta t=10^{-2}$, iters=10
$1/16$ 1.01041e-01 0.00 6.10036e-03 0.00 2.71376e-01 0.00 2.39870e-01 0.00
$1/32$ 2.54196e-02 1.99 1.51238e-03 2.01 1.18975e-01 1.19 1.11050e-01 1.11
$1/64$ 6.36474e-03 2.00 3.75044e-04 2.01 5.60370e-02 1.09 5.43245e-02 1.03
$1/128$ 1.59163e-03 2.00 9.15599e-05 2.03 2.72366e-02 1.04 2.70131e-02 1.01
Table 5: Example 1: The errors and convergence orders with decreasing mesh sizes of coupled and iterative decoupling algorithms. $\bm{\Theta}=10^{-6}\bm{I}$
Alg.1, $\Delta t=10^{-3}$
$1/16$ 1.01039e-01 0.00 6.07933e-03 0.00 2.35294e-01 0.00 2.66261e-01 0.00
$1/32$ 2.54193e-02 1.99 1.50919e-03 2.01 1.10485e-01 1.09 1.18339e-01 1.17
$1/64$ 6.36485e-03 2.00 3.76523e-04 2.00 5.42944e-02 1.02 5.62194e-02 1.07
$1/128$ 1.59182e-03 2.00 9.38336e-05 2.00 2.70270e-02 1.01 2.74790e-02 1.03
Alg.2, $\Delta t=5\times 10^{-3}$, iters=5
$1/16$ 1.01040e-01 0.00 6.08908e-03 0.00 2.37202e-01 0.00 2.68502e-01 0.00
$1/32$ 2.54196e-02 1.99 1.51189e-03 2.01 1.10762e-01 1.10 1.18668e-01 1.18
$1/64$ 6.36496e-03 2.00 3.77393e-04 2.00 5.43382e-02 1.03 5.62122e-02 1.08
$1/128$ 1.59189e-03 2.00 9.43212e-05 2.00 2.70369e-02 1.01 2.74213e-02 1.04
Alg.2, $\Delta t=10^{-2}$, iters=10
$1/16$ 1.01041e-01 0.00 6.10036e-03 0.00 2.39870e-01 0.00 2.71376e-01 0.00
$1/32$ 2.54196e-02 1.99 1.51238e-03 2.01 1.11050e-01 1.11 1.18975e-01 1.19
$1/64$ 6.36474e-03 2.00 3.75044e-04 2.01 5.43245e-02 1.03 5.60370e-02 1.09
$1/128$ 1.59163e-03 2.00 9.15599e-05 2.03 2.70131e-02 1.01 2.72366e-02 1.04
Table 6: Example 1: The errors and convergence orders with decreasing mesh sizes of coupled and iterative decoupling algorithms. $\bm{K}=10^{-6}\bm{I},\bm{\Theta}=10^{-6}\bm{I}$
Alg.1, $\Delta t=10^{-3}$
$1/16$ 1.01060e-01 0.00 6.30251e-03 0.00 3.13011e-01 0.00 3.13011e-01 0.00
$1/32$ 2.54251e-02 1.99 1.57062e-03 2.00 1.32705e-01 1.24 1.32705e-01 1.24
$1/64$ 6.36635e-03 2.00 3.92445e-04 2.00 6.01621e-02 1.14 6.01621e-02 1.14
$1/128$ 1.59219e-03 2.00 9.77413e-05 2.01 2.84617e-02 1.08 2.84617e-02 1.08
Alg.2, $\Delta t=5\times 10^{-3}$, iters=5
$1/16$ 1.01060e-01 0.00 6.29942e-03 0.00 3.12563e-01 0.00 3.12563e-01 0.00
$1/32$ 2.54252e-02 1.99 1.57117e-03 2.00 1.32870e-01 1.23 1.32870e-01 1.23
$1/64$ 6.36653e-03 2.00 3.94108e-04 2.00 6.05260e-02 1.13 6.05260e-02 1.13
$1/128$ 1.59241e-03 2.00 9.99148e-05 1.98 2.89124e-02 1.07 2.89124e-02 1.07
Alg.2, $\Delta t=10^{-2}$, iters=10
$1/16$ 1.01060e-01 0.00 6.29765e-03 0.00 3.12246e-01 0.00 3.12246e-01 0.00
$1/32$ 2.54246e-02 1.99 1.56551e-03 2.01 1.31875e-01 1.24 1.31875e-01 1.24
$1/64$ 6.36589e-03 2.00 3.87375e-04 2.01 5.93113e-02 1.15 5.93113e-02 1.15
$1/128$ 1.59179e-03 2.00 9.32757e-05 2.05 2.77070e-02 1.10 2.77070e-02 1.10
@@ 5.3 Tests for the parameter $a_{0},b_{0},c_{0}$
Table 7: Example 1: The errors and convergence orders with decreasing mesh sizes of coupled and iterative decoupling algorithms. $a_{0}=b_{0}=c_{0}=0$
@@
$1/16$ 1.01017e-01 0.00 5.86499e-03 0.00 2.18610e-01 0.00 2.18610e-01 0.00
$1/32$ 2.54138e-02 1.99 1.45400e-03 2.01 1.08332e-01 1.01 1.08332e-01 1.01
$1/64$ 6.36347e-03 2.00 3.62641e-04 2.00 5.40235e-02 1.00 5.40235e-02 1.00
$1/128$ 1.59148e-03 2.00 9.04325e-05 2.00 2.69935e-02 1.00 2.69935e-02 1.00
Alg.2, $\Delta t=5\times 10^{-3}$, iters=5
$1/16$ 1.01050e-01 0.00 6.19651e-03 0.00 2.48520e-01 0.00 2.48520e-01 0.00
$1/32$ 2.54262e-02 1.99 1.57535e-03 1.98 1.14200e-01 1.12 1.14200e-01 1.12
$1/64$ 6.37219e-03 2.00 4.59498e-04 1.78 5.61495e-02 1.02 5.61495e-02 1.02
$1/128$ 1.60655e-03 1.99 2.39721e-04 0.94 2.88472e-02 0.96 2.88472e-02 0.96
Alg.2, $\Delta t=10^{-2}$, iters=10
$1/16$ 1.01126e-01 0.00 6.95719e-03 0.00 3.08253e-01 0.00 3.08253e-01 0.00
$1/32$ 2.54412e-02 1.99 1.73121e-03 2.01 1.21121e-01 1.35 1.21121e-01 1.35
$1/64$ 6.36954e-03 2.00 4.24246e-04 2.03 5.55013e-02 1.13 5.55013e-02 1.13
$1/128$ 1.59228e-03 2.00 9.84554e-05 2.11 2.70932e-02 1.03 2.70932e-02 1.03
5.4 Tests for CPU time
In this subsection, we examine the computational cost, which is represented by the indicator of CPU times (in seconds). We maintain the same set of parameters: $\nu=0.3$, $E=1$ ,$a_{0}=c_{0}=0.2$, $b_{0}=0.1$, $\alpha=\beta=0.1$, and $\bm{K}=\bm{\Theta}=0.1\bm{I}$. The terminal time is set to $\tau=1$ to make the runtime more noticeable. We have presented Table LABEL:tab:_time_compare, which provides a comparison of CPU times across three scenarios: Algorithm 1 with a time step $\Delta t=10^{-3}$, Algorithm 2 with a time step $\Delta t=5\times 10^{-3}$ and 5 iterations per time step, and Algorithm 1 with a time step $\Delta t=10^{-2}$ and 10 iterations per step. We consider two different mesh sizes, $h=1/20$ and $h=1/40$, respectively. Table LABEL:tab:_time_compare reveals that when the error approaches a similar level, the computational cost of Algorithm 1 exceeds that of Algorithm 2. Additionally, as the time step increases, the computational cost decreases.
Table 8: CPU time(s) when $\nu=0.3,E=1,a_{0}=c_{0}=0.2,b_{0}=0.1,\alpha=\beta=0.1,\bm{K}=\bm{\Theta}=0.1\bm{I}$
iters $h$ $\Delta t$ $\|\bm{u}\|_{H^{1}}$ $\|\xi\|_{L^{2}}$ $\|p\|_{H^{1}}$ $\|T\|_{H^{1}}$ CPU time(s)
Algorithm 1 - $1/20$ $10^{-3}$ 2.40947e-02 1.38254e-03 6.40944e-02 6.40944e-02 421.69
Algorithm 2 $5$ $1/20$ $5\times 10^{-3}$ 2.40947e-02 1.38136e-03 6.40931e-02 6.40931e-02 304.11
Algorithm 2 $01$/201201/201 / 20 $10^{-2}$ 2.40947e-02 1.38086e-03 6.40939e-02 6.40939e-02 153.95
Algorithm 1 - $1/40$ $10^{-3}$ 6.04800e-03 3.43627e-04 3.20810e-02 3.20810e-02 1141.04
Algorithm 2 $5$ $1/40$ $5\times 10^{-3}$ 6.04799e-03 3.42506e-04 3.20809e-02 3.20809e-02 858.47
Algorithm 2 $01$/401401/401 / 40 $10^{-2}$ 6.04799e-03 3.42132e-04 3.20836e-02 3.20836e-02 427.85
5.5 Tests for a large terminal time
Finally, in this subsection, we consider the performance of the iterative decoupled algorithm at a large terminal time $\tau=1$. We directly test the extreme situation where $\bm{K}=\bm{\Theta}=10^{-6}\bm{I}$ and maintain the other parameters as $\nu=0.3$, $E=1$, $a_{0}=c_{0}=0.2$, $b_{0}=0.1$ and $\alpha=\beta=1$. In Table LABEL:tab:_tau=1, we use the same time step for both Algorithm 1 (Alg. 1) and Algorithm 2 (Alg. 2). For Algorithm 2, we consider two iteration numbers: iters = 5 and iters = 10. We found that as the number of $iters$ increases from 5 to 10, the errors of Algorithm 2 become closer to those of Algorithm 1, which consolidates Theorem 4.1, which describes that the solutions of Algorithm 2 converge to the solutions of Algorithm 1 at each time step.
Table 9: Example 1: The errors and convergence orders with decreasing mesh sizes of coupled and iterative decoupling algorithms. $\tau=1$, $\bm{K}=\bm{\Theta}=10^{-6}\bm{I}$
Alg.1, $\Delta t=2.5\times 10^{-3}$
$1/8$ 1.46881e-01 0.00 1.28807e-02 0.00 6.85628e-01 0.00 6.85628e-01 0.00
$1/16$ 3.76413e-02 1.96 3.16363e-03 2.03 2.22414e-01 1.62 2.22414e-01 1.62
$1/32$ 9.45795e-03 1.99 6.86969e-04 2.20 6.46414e-02 1.78 6.46414e-02 1.78
$1/64$ 2.36534e-03 2.00 1.46263e-04 2.23 2.20416e-02 1.55 2.20416e-02 1.55
Alg.2, $\Delta t=2.5\times 10^{-3}$, iters=5
$1/8$ 1.46934e-01 0.00 1.32454e-02 0.00 7.24589e-01 0.00 7.24589e-01 0.00
$1/16$ 3.77088e-02 1.96 3.63632e-03 1.86 2.78004e-01 1.38 2.78004e-01 1.38
$1/32$ 9.54308e-03 1.98 1.25329e-03 1.54 1.41638e-01 0.97 1.41638e-01 0.97
$1/64$ 2.50093e-03 1.93 7.31934e-04 0.78 1.13157e-01 0.32 1.13157e-01 0.32
Alg.2, $\Delta t=2.5\times 10^{-3}$, iters=10
$1/8$ 1.46881e-01 0.00 1.28808e-02 0.00 6.85633e-01 0.00 6.85633e-01 0.00
$1/16$ 3.76413e-02 1.96 3.16367e-03 2.03 2.22421e-01 1.62 2.22421e-01 1.62
$1/32$ 9.45796e-03 1.99 6.87016e-04 2.20 6.46508e-02 1.78 6.46508e-02 1.78
$1/64$ 2.36534e-03 2.00 1.46274e-04 2.23 2.20355e-02 1.55 2.20355e-02 1.55

### v3

@@ 5.2 Tests for the parameters $K$ and $\Theta$
Table 3: The errors and convergence orders with decreasing mesh sizes $\bm{K}=10^{-6}\bm{I}$.
$h$ $\|\bm{u}\|_{H^{1}}$ rate $\|\xi\|_{L^{2}}$ rate $\|p\|_{H^{1}}$ rate $\|T\|_{H^{1}}$ rate
Alg.1,
$\Delta t=10^{-3}$
$1/16$ 1.00615e-01 0.00 6.07391e-03 0.00 2.48018e-01 0.00 2.31172e-01 0.00
$1/32$ 2.53670e-02 1.99 1.50130e-03 2.02 1.13982e-01 1.12 1.09927e-01 1.07
$1/64$ 6.35858e-03 2.00 3.74103e-04 2.00 5.51891e-02 1.05 5.42244e-02 1.02
$1/128$ 1.59110e-03 2.00 9.32365e-05 2.00 2.72419e-02 1.02 2.70188e-02 1.00
Alg.2,
$\Delta t=5\times 10^{-3}$,
iters=5
$1/16$ 1.00615e-01 0.00 6.07747e-03 0.00 2.49194e-01 0.00 2.32137e-01 0.00
$1/32$ 2.53671e-02 1.99 1.50226e-03 2.02 1.14131e-01 1.13 1.10059e-01 1.08
$1/64$ 6.35859e-03 2.00 3.74397e-04 2.00 5.51695e-02 1.05 5.42417e-02 1.02
$1/128$ 1.59111e-03 2.00 9.33786e-05 2.00 2.72016e-02 1.02 2.70214e-02 1.01
Alg.2,
$\Delta t=10^{-2}$,
iters=10
$1/16$ 1.00616e-01 0.00 6.08155e-03 0.00 2.50703e-01 0.00 2.33492e-01 0.00
$1/32$ 2.53669e-02 1.99 1.50131e-03 2.02 1.14277e-01 1.13 1.10188e-01 1.08
$1/64$ 6.35831e-03 2.00 3.72183e-04 2.01 5.50831e-02 1.05 5.42265e-02 1.02
$1/128$ 1.59083e-03 2.00 9.11603e-05 2.03 2.71163e-02 1.02 2.70061e-02 1.01
Table 4: The errors and convergence orders with decreasing mesh sizes. $\bm{\Theta}=10^{-6}\bm{I}$.
$h$ $\|\bm{u}\|_{H^{1}}$ rate $\|\xi\|_{L^{2}}$ rate $\|p\|_{H^{1}}$ rate $\|T\|_{H^{1}}$ rate
Alg.1,
$\Delta t=10^{-3}$
$1/16$ 1.00615e-01 0.00 6.07391e-03 0.00 2.31172e-01 0.00 2.48018e-01 0.00
$1/32$ 2.53670e-02 1.99 1.50130e-03 2.02 1.09927e-01 1.07 1.13982e-01 1.12
$1/64$ 6.35858e-03 2.00 3.74103e-04 2.00 5.42244e-02 1.02 5.51891e-02 1.05
$1/128$ 1.59110e-03 2.00 9.32365e-05 2.00 2.70188e-02 1.00 2.72419e-02 1.02
Alg.2,
$\Delta t=5\times 10^{-3}$,
iters=5
$1/16$ 1.00615e-01 0.00 6.07747e-03 0.00 2.32137e-01 0.00 2.49194e-01 0.00
$1/32$ 2.53671e-02 1.99 1.50226e-03 2.02 1.10059e-01 1.08 1.14131e-01 1.13
$1/64$ 6.35859e-03 2.00 3.74397e-04 2.00 5.42417e-02 1.02 5.51695e-02 1.05
$1/128$ 1.59111e-03 2.00 9.33786e-05 2.00 2.70214e-02 1.01 2.72016e-02 1.02
Alg.2,
$\Delta t=10^{-2}$,
iters=10
$1/16$ 1.00616e-01 0.00 6.08155e-03 0.00 2.33492e-01 0.00 2.50703e-01 0.00
$1/32$ 2.53669e-02 1.99 1.50131e-03 2.02 1.10188e-01 1.08 1.14277e-01 1.13
$1/64$ 6.35831e-03 2.00 3.72183e-04 2.01 5.42265e-02 1.02 5.50831e-02 1.05
$1/128$ 1.59083e-03 2.00 9.11603e-05 2.03 2.70061e-02 1.01 2.71163e-02 1.02
Table 5: The errors and convergence orders with decreasing mesh sizes. $\bm{K}=10^{-6}\bm{I},\bm{\Theta}=10^{-6}\bm{I}$.
$h$ $\|\bm{u}\|_{H^{1}}$ rate $\|\xi\|_{L^{2}}$ rate $\|p\|_{H^{1}}$ rate $\|T\|_{H^{1}}$ rate
Alg.1,
$\Delta t=10^{-3}$
$1/16$ 1.00629e-01 0.00 6.18582e-03 0.00 2.73973e-01 0.00 2.73973e-01 0.00
$1/32$ 2.53705e-02 1.99 1.53143e-03 2.01 1.21482e-01 1.17 1.21482e-01 1.17
$1/64$ 6.35948e-03 2.00 3.81830e-04 2.00 5.71664e-02 1.09 5.71664e-02 1.09
$1/128$ 1.59132e-03 2.00 9.51228e-05 2.01 2.77233e-02 1.04 2.77233e-02 1.04
Alg.2,
$\Delta t=5\times 10^{-3}$,
iters=5
$1/16$ 1.00628e-01 0.00 6.18391e-03 0.00 2.73713e-01 0.00 2.73713e-01 0.00
$1/32$ 2.53705e-02 1.99 1.53140e-03 2.01 1.21507e-01 1.17 1.21507e-01 1.17
$1/64$ 6.35951e-03 2.00 3.82333e-04 2.00 5.72693e-02 1.09 5.72693e-02 1.09
$1/128$ 1.59137e-03 2.00 9.58070e-05 2.00 2.78597e-02 1.04 2.78597e-02 1.04
Alg.2,
$\Delta t=10^{-2}$,
iters=10
$1/16$ 1.00628e-01 0.00 6.18228e-03 0.00 2.73490e-01 0.00 2.73490e-01 0.00
$1/32$ 2.53701e-02 1.99 1.52778e-03 2.02 1.21011e-01 1.18 1.21011e-01 1.18
$1/64$ 6.35902e-03 2.00 3.78244e-04 2.01 5.67190e-02 1.09 5.67190e-02 1.09
$1/128$ 1.59093e-03 2.00 9.19929e-05 2.04 2.73453e-02 1.05 2.73453e-02 1.05
@@ 5.3 Tests for the parameter $a_{0},b_{0},c_{0}$
Table 6: The errors and convergence orders with decreasing mesh sizes. $a_{0}=b_{0}=c_{0}=0$.
Alg.1,
$\Delta t=10^{-3}$
$1/16$ 1.00716e-01 0.00 6.74538e-03 0.00 2.60803e-01 0.00 2.60803e-01 0.00
$1/32$ 2.53932e-02 1.99 1.67613e-03 2.01 1.14106e-01 1.19 1.14106e-01 1.19
$1/64$ 6.36502e-03 2.00 4.17182e-04 2.01 5.47500e-02 1.06 5.47500e-02 1.06
$1/128$ 1.59255e-03 2.00 1.02942e-04 2.02 2.70774e-02 1.02 2.70774e-02 1.02
Alg.2,
$\Delta t=5\times 10^{-3}$,
iters=5
$1/16$ 1.00807e-01 0.00 7.31400e-03 0.00 2.91681e-01 0.00 2.91681e-01 0.00
$1/32$ 2.54374e-02 1.99 1.95023e-03 1.91 1.22342e-01 1.25 1.22342e-01 1.25
$1/64$ 6.40628e-03 1.99 6.41686e-04 1.60 5.85438e-02 1.06 5.85438e-02 1.06
$1/128$ 1.66723e-03 1.94 3.49812e-04 0.88 3.04263e-02 0.94 3.04263e-02 0.94
Alg.2,
$\Delta t=10^{-2}$,
iters=10
$1/16$ 1.00916e-01 0.00 7.96738e-03 0.00 3.30774e-01 0.00 3.30774e-01 0.00
$1/32$ 2.54412e-02 1.99 1.97495e-03 2.01 1.24587e-01 1.41 1.24587e-01 1.41
$1/64$ 6.37477e-03 2.00 4.79389e-04 2.04 5.59219e-02 1.16 5.59219e-02 1.16
$1/128$ 1.59309e-03 2.00 1.06986e-04 2.16 2.71232e-02 1.04 2.71232e-02 1.04
@@
5.4 Tests with a large terminal time $\tau$
In this subsection, we evaluate the performance of the proposed algorithms for a relatively long terminal time $\tau=1$. Owing to the extended simulation duration, all algorithms require a greater number of time steps to complete the computation. The parameters are set as follows: $\nu=0.3$, $E=1$, $a_{0}=c_{0}=0.2$, $b_{0}=0.1$, $\alpha=\beta=0.1$, and $\bm{K}=\bm{\Theta}=0.1I$.
Table LABEL:tab:_same_Dt_of_Alg1_and_Alg2 presents a comparison of Alg. 1 and Alg. 2 under different time-step sizes, specifically $\Delta t=10^{-3}$, $\Delta t=5\times 10^{-3}$, and $\Delta t=10^{-2}$. In all these cases, the spatial discretization error remains dominant. The numerical results show that both the coupled and decoupled algorithms yield comparable accuracy. As expected, reducing the time-step size leads to improved accuracy for both algorithms. Furthermore, for fixed time-step sizes, increasing the number of iterations in Alg. 2 from 5 to 10 results in errors that more closely align with those of Alg. 1, thereby confirming the theoretical convergence behavior stated in Theorem 4.1.
As in the comparisons presented in Table LABEL:tab:_K_Theta_0.1_nu_0.3, we examine Alg. 1 using a time-step size of $\Delta t=1.0\times 10^{-3}$, and Alg. 2 using $\Delta t=5\times 10^{-3}$ with 5 iterations or $\Delta t=10^{-2}$ with 10 iterations. The results show that the numerical errors for both algorithms are comparable. These observations confirm that the iterative decoupled algorithm remains accurate and robust even when applied to simulations with large terminal times. Moreover, consistent with the findings in Table LABEL:tab:_K_Theta_0.1_nu_0.3, the decoupled algorithm offers improved computational efficiency.
Table 7: The errors and convergence orders of different algorithms. $\tau=1$.
@@ 5.4 Tests with a large terminal time $\tau$
Alg.1,$\Delta t=10^{-2}$
$1/8$ 1.45210e-01 0.00 9.12749e-03 0.00 1.58002e-01 0.00 1.58002e-01 0.00
$1/16$ 3.73731e-02 1.96 2.16086e-03 2.08 7.99221e-02 0.98 7.99221e-02 0.98
$1/32$ 9.42251e-03 1.99 5.33182e-04 2.02 4.00855e-02 1.00 4.00855e-02 1.00
$1/64$ 2.36208e-03 2.00 1.34243e-04 1.99 2.00727e-02 1.00 2.00727e-02 1.00
Alg.2,$\Delta t=10^{-2}$,iters=5
$1/8$ 1.45210e-01 0.00 9.12737e-03 0.00 1.58000e-01 0.00 1.58000e-01 0.00
$1/16$ 3.73731e-02 1.96 2.16072e-03 2.08 7.99198e-02 0.98 7.99198e-02 0.98
$1/32$ 9.42248e-03 1.99 5.32938e-04 2.02 4.00813e-02 1.00 4.00813e-02 1.00
$1/64$ 2.36200e-03 2.00 1.33584e-04 2.00 2.00645e-02 1.00 2.00645e-02 1.00
Alg.2,$\Delta t=10^{-2}$,iters=10
$1/8$ 1.45210e-01 0.00 9.12749e-03 0.00 1.58002e-01 0.00 1.58002e-01 0.00
$1/16$ 3.73731e-02 1.96 2.16086e-03 2.08 7.99221e-02 0.98 7.99221e-02 0.98
$1/32$ 9.42251e-03 1.99 5.33182e-04 2.02 4.00855e-02 1.00 4.00855e-02 1.00
$1/64$ 2.36208e-03 2.00 1.34243e-04 1.99 2.00727e-02 1.00 2.00727e-02 1.00
Alg.1,$\Delta t=5\times 10^{-3}$
$1/8$ 1.45210e-01 0.00 9.12771e-03 0.00 1.58002e-01 0.00 1.58002e-01 0.00
$1/16$ 3.73731e-02 1.96 2.16106e-03 2.08 7.99192e-02 0.98 7.99192e-02 0.98
$1/32$ 9.42250e-03 1.99 5.33110e-04 2.02 4.00783e-02 1.00 4.00783e-02 1.00
$1/64$ 2.36193e-03 2.00 1.33096e-04 2.00 2.00576e-02 1.00 2.00576e-02 1.00
Alg.2,$\Delta t=5\times 10^{-3}$,iters=5
$1/8$ 1.45210e-01 0.00 9.12758e-03 0.00 1.58001e-01 0.00 1.58001e-01 0.00
$1/16$ 3.73731e-02 1.96 2.16092e-03 2.08 7.99180e-02 0.98 7.99180e-02 0.98
$1/32$ 9.42248e-03 1.99 5.32940e-04 2.02 4.00764e-02 1.00 4.00764e-02 1.00
$1/64$ 2.36189e-03 2.00 1.32775e-04 2.00 2.00541e-02 1.00 2.00541e-02 1.00
Alg.2,$\Delta t=5\times 10^{-3}$,iters=10
$1/8$ 1.45210e-01 0.00 9.12771e-03 0.00 1.58002e-01 0.00 1.58002e-01 0.00
$1/16$ 3.73731e-02 1.96 2.16106e-03 2.08 7.99192e-02 0.98 7.99192e-02 0.98
$1/32$ 9.42250e-03 1.99 5.33110e-04 2.02 4.00783e-02 1.00 4.00783e-02 1.00
$1/64$ 2.36193e-03 2.00 1.33096e-04 2.00 2.00576e-02 1.00 2.00576e-02 1.00
@@
$1/8$ 1.45210e-01 0.00 9.12790e-03 0.00 1.58003e-01 0.00 1.58003e-01 0.00
$1/16$ 3.73731e-02 1.96 2.16126e-03 2.08 7.99186e-02 0.98 7.99186e-02 0.98
$1/32$ 9.42251e-03 1.99 5.33223e-04 2.02 4.00762e-02 1.00 4.00762e-02 1.00
$1/64$ 2.36189e-03 2.00 1.32863e-04 2.00 2.00529e-02 1.00 2.00529e-02 1.00
Alg.2,$\Delta t=10^{-3}$,iters=5
$1/8$ 1.45210e-01 0.00 9.12776e-03 0.00 1.58002e-01 0.00 1.58002e-01 0.00
$1/16$ 3.73731e-02 1.96 2.16113e-03 2.08 7.99189e-02 0.98 7.99189e-02 0.98
$1/32$ 9.42250e-03 1.99 5.33150e-04 2.02 4.00774e-02 1.00 4.00774e-02 1.00
$1/64$ 2.36191e-03 2.00 1.32987e-04 2.00 2.00556e-02 1.00 2.00556e-02 1.00
Alg.2,$\Delta t=10^{-3}$,iters=10
$1/8$ 1.45210e-01 0.00 9.12790e-03 0.00 1.58003e-01 0.00 1.58003e-01 0.00
$1/16$ 3.73731e-02 1.96 2.16126e-03 2.08 7.99186e-02 0.98 7.99186e-02 0.98
$1/32$ 9.42251e-03 1.99 5.33223e-04 2.02 4.00762e-02 1.00 4.00762e-02 1.00
$1/64$ 2.36189e-03 2.00 1.32863e-04 2.00 2.00529e-02 1.00 2.00529e-02 1.00